# Topological zero modes and bounded modes at smooth domain walls: exact solutions and dualities

### v1

@@ Introduction
[…]
Table 5: Duality between the zero modes of the modified Jackiw-Rebbi equation and the bounded modes of a Schrödinger equation under the transformation $v(x)^{2}+sv^{\prime}(x)+m(x)=V(x)-E$ on the infinite interval $-\infty<x<\infty$ and on the semi-infinite interval $0\leq x<\infty$. Notice that $\widetilde{W}_{L,R}=1$ mandates $|W_{L,R}|=1$ assuming $v_{L,R}\neq 0$ and that $\widetilde{W}_{L,R}=0$ mandates $|W_{L,R}|=0$ assuming $|m_{L,R}|>v_{L,R}^{2}$ Moreover one has $v^{2}_{L,R}+m_{L,R}=V_{L,R}-E$ in all cases. Hence, there is a one-to-one correspondence between bounded modes and zero modes and between the oscillations of the bounded modes and damped oscillatons of the zero modes.
Hence, surprisingly, the zero modes of the modified Jackiw-Rebbi equation describing are directly related to the bounded modes of a Schrödinger equation of a particle scattering off a potential. However, there are some important differences: There are now no topological invariant, since the Hamiltonian is now topologically trivial for all parameters, no localized zero modes at the phase boundary, and consequently no bulk-boundary correspondence. However, we can recover a sort of bulk-boundary correspondence governing the existence of zero modes at phase boundaries, if the role of the topological invariant is now played by quantity $\widetilde{W}$. If the quantity $\widetilde{W}$ changes (i.e., if $V-E$ changes sign) in correspondence of the smooth domain wall, there exists a localized Shockley-like mode which exponentially decays (or exhibits exponentially damped oscillations) on the side of the domain wall with $\widetilde{W}=0$ (the side with $V-E>0$). If the $\widetilde{W}$ does not change at the domain wall, there can exist only wavelike modes, or no modes at all. Notice that this also forbid the existence of Shockley-like modes at the left end of the interval $0\leq x<\infty$ if $\widetilde{W}_{R}=0$, assigning $\widetilde{W}_{L}=0$ to the "vacuum", i.e., by taking $V(x)\to\infty$ for $x<0$ (i.e., $V-E>0$). This identification of $\widetilde{W}$ with a topological invariant and the consequent interpretation of localized Shockley-like modes in terms of a bulk-boundary correspondence is not based on topological considerations must be taken with a grain of salt.
Similarly to the case of the zero modes solutions of the modified Jackiw-Rebbi equation, to gain a physical intuition on the wavefunction of the bounded modes at short distances $x\lesssim w$, one can analize the spatial dependence of $V(x)-E$ and define the quantity
@@
which generalizes $\widetilde{W}_{L,R}$ defined in Eq. 22 to the whole interval $-\infty<x<\infty$. We have seen that the character of the bounded modes on the interval $-\infty<x<\infty$ at long distances on each side of the smooth boundary depends on the value of $\widetilde{W}_{L,R}$: If $\widetilde{W}_{L,R}=1$, the mode has a wavelike character, converging to a constant value or oscillating with no decay, while if $\widetilde{W}_{L,R}=0$ the mode has a Shockley-like character with exponentially decay.
@@
Figure 3: Bounded modes on the interval $-\infty<x<\infty$ with a smooth domain wall given by the field $V(x)$ and $E=0$.Upper panels show the wavefunction and lower panes show the field $V(x)$. Different columns correspond via the duality transformation to the columns in Fig. 2, by taking $V_{0}=(v_{0}^{2}+m_{0})$, $V_{1}=(2v_{0}v_{1}+(s/w)v_{1}+m_{1})$, and $V_{2}=(v_{1}^{2}-(s/w)v_{1}+m_{2})$ with $s=1$. This corresponds to: (a) single Shockley-like mode with exponential decay on the left side and oscillations on the right side of a smooth domain wall with S-shaped field $V(x)$ [corresponding to Fig. 2(b)]; (b) two wavelike modes with oscillations on both sides of a smooth domain wall with symmetric Pöschl–Teller field $V(x)$ [corresponding to Fig. 2(c) and taking the two linear combinations $\widetilde{\varphi}_{1}\pm\widetilde{\varphi}_{2}$ giving real wavefunctions]; (c) single Shockley-like mode with exponential decay on the left side and oscillations on the right side of a smooth domain wall with asymmetric Pöschl–Teller $V(x)$ [corresponding to Fig. 2(e)]; (d) two wavelike modes with oscillations on both sides of a smooth domain wall with symmetric Pöschl–Teller field $V(x)$ [corresponding to Fig. 2(f) and taking the two linear combinations $\widetilde{\varphi}_{1}\pm\widetilde{\varphi}_{2}$].
@@
Figure 4 show bounded modes on the interval $-\infty<x<\infty$ for different choices of the parameters $V_{0,1,2}$ and with $E=0$ plotted using Eqs. 21 and 5. Figure 4 shows the case where the width of the smooth domain wall $w$ becomes larger. We choose the parameters such $V_{0,1,2}$ to correspond to the parameters $m_{0,1,2}$ and $v_{1,2}$ via the duality transformations, i.e., taking $V_{0}=(v_{0}^{2}+m_{0})$, $V_{1}=(2v_{0}v_{1}+(s/w)v_{1}+m_{1})$, and $V_{2}=(v_{1}^{2}-(s/w)v_{1}+m_{2})$ with $m_{0,1,2}$ and $v_{0,1}$ as in Fig. 2(b), Fig. 2(c),Fig. 2(e), Fig. 2(f), to obtain Fig. 4(a), Fig. 4(b), Fig. 4(c), Fig. 4(d), respectively, and as in Fig. 2(b), Fig. 2(c), Fig. 2(e), Fig. 2(f), to obtain Fig. 4(a), Fig. 4(b),Fig. 4(c), Fig. 4(d), respectively. Analogously to the case of zero modes, the local features of the bounded modes become prominent when the width of the smooth domain wall increases and when the field around the domain wall is large. This is especially visible in Fig. 4(c), where the crossover between oscillations and decaying behavior occurs far away from the center of the domain wall, in correspondence with a point where the quantity $\widetilde{W}(x)$ changes, i.e., where the field $V(x)$ changes its sign.

### v2

@@ Introduction
1 wavelike mode $\sum A_{i}\widetilde{\varphi}_{i}$ $\overleftrightarrow{\qquad\qquad\qquad\qquad\qquad\qquad\qquad}$ 1 zero mode $\sum A_{i}\varphi_{i}^{s}$
$V_{R}=E\Rightarrow$ const. at +$\infty$+ ∞ $\overleftrightarrow{\qquad\qquad\qquad\qquad\qquad\qquad\qquad}$ $v_{R}^{2}+m_{R}=0\Rightarrow$ exp. decay at +$\infty$+ ∞
$V_{R}<E\Rightarrow$ oscill. at +$\infty$+ ∞ $\overleftrightarrow{\qquad\qquad\qquad\qquad\qquad\qquad\qquad}$ $v_{R}^{2}+m_{R}<0\Rightarrow$ exp. damped oscill. at +$\infty$+ ∞
Table 5: Duality between the zero modes of the modified Jackiw-Rebbi equation and the bounded modes of a Schrödinger equation under the transformation $v(x)^{2}+sv^{\prime}(x)+m(x)=V(x)-E$ on the infinite interval $-\infty<x<\infty$ and on the semi-infinite interval $0\leq x<\infty$. Notice that $\widetilde{W}_{L,R}=1$ mandates $|W_{L,R}|=1$ assuming $v_{L,R}\neq 0$ and that $\widetilde{W}_{L,R}=0$ mandates $|W_{L,R}|=0$ assuming $|m_{L,R}|>v_{L,R}^{2}$ Moreover one has $v^{2}_{L,R}+m_{L,R}=V_{L,R}-E$ in all cases. Hence, there is a one-to-one correspondence between bounded modes and zero modes and between the oscillations of the bounded modes and damped oscillations of the zero modes.
Hence, surprisingly, the zero modes of the modified Jackiw-Rebbi equation are directly related to the bounded modes of a Schrödinger equation of a particle scattering off a potential. However, there are some important differences: There are now no topological invariant, since the Hamiltonian is now topologically trivial for all parameters, no localized zero modes at the phase boundary, and consequently no bulk-boundary correspondence. However, we can recover a sort of bulk-boundary correspondence governing the existence of zero modes at phase boundaries, if the role of the topological invariant is now played by quantity $\widetilde{W}$. If the quantity $\widetilde{W}$ changes (i.e., if $V-E$ changes sign) in correspondence of the smooth domain wall, there exists a localized Shockley-like mode which exponentially decays (or exhibits exponentially damped oscillations) on the side of the domain wall with $\widetilde{W}=0$ (the side with $V-E>0$). If the $\widetilde{W}$ does not change at the domain wall, there can exist only wavelike modes, or no modes at all. Notice that this also forbid the existence of Shockley-like modes at the left end of the interval $0\leq x<\infty$ if $\widetilde{W}_{R}=0$, assigning $\widetilde{W}_{L}=0$ to the "vacuum", i.e., by taking $V(x)\to\infty$ for $x<0$ (i.e., $V-E>0$). This identification of $\widetilde{W}$ with a topological invariant and the consequent interpretation of localized Shockley-like modes in terms of a bulk-boundary correspondence is not based on topological considerations must be taken with a grain of salt.
Similarly to the case of the zero modes solutions of the modified Jackiw-Rebbi equation, to gain a physical intuition on the wavefunction of the bounded modes at short distances $x\lesssim w$, one can analyze the spatial dependence of $V(x)-E$ and define the quantity
@@
which generalizes $\widetilde{W}_{L,R}$ defined in Eq. 24 to the whole interval $-\infty<x<\infty$. We have seen that the character of the bounded modes on the interval $-\infty<x<\infty$ at long distances on each side of the smooth boundary depends on the value of $\widetilde{W}_{L,R}$: If $\widetilde{W}_{L,R}=1$, the mode has a wavelike character, converging to a constant value or oscillating with no decay, while if $\widetilde{W}_{L,R}=0$ the mode has a Shockley-like character with exponential decay.
@@
Figure 3: Bounded modes on the interval $-\infty<x<\infty$ with a smooth domain wall given by the field $V(x)$ and $E=0$. Upper panels show the wavefunction and lower panels show the field $V(x)$. Different columns correspond via the duality transformation to the columns in Fig. 2, by taking $V_{0}=(v_{0}^{2}+m_{0})$, $V_{1}=(2v_{0}v_{1}+(s/w)v_{1}+m_{1})$, and $V_{2}=(v_{1}^{2}-(s/w)v_{1}+m_{2})$ with $s=1$. This corresponds to: (a) single Shockley-like mode with exponential decay on the left side and oscillations on the right side of a smooth domain wall with S-shaped field $V(x)$ [corresponding to Fig. 2(b)]; (b) two wavelike modes with oscillations on both sides of a smooth domain wall with symmetric Pöschl–Teller field $V(x)$ [corresponding to Fig. 2(c) and taking the two linear combinations $\widetilde{\varphi}_{1}\pm\widetilde{\varphi}_{2}$ giving real wavefunctions]; (c) single Shockley-like mode with exponential decay on the left side and oscillations on the right side of a smooth domain wall with asymmetric Pöschl–Teller $V(x)$ [corresponding to Fig. 2(e)]; (d) two wavelike modes with oscillations on both sides of a smooth domain wall with symmetric Pöschl–Teller field $V(x)$ [corresponding to Fig. 2(f) and taking the two linear combinations $\widetilde{\varphi}_{1}\pm\widetilde{\varphi}_{2}$].
@@
Figure 4 show bounded modes on the interval $-\infty<x<\infty$ for different choices of the parameters $V_{0,1,2}$ and with $E=0$ plotted using Eqs. 23 and 7. Figure 4 shows the case where the width of the smooth domain wall $w$ becomes larger. We choose the parameters such $V_{0,1,2}$ to correspond to the parameters $m_{0,1,2}$ and $v_{1,2}$ via the duality transformations, i.e., taking $V_{0}=(v_{0}^{2}+m_{0})$, $V_{1}=(2v_{0}v_{1}+(s/w)v_{1}+m_{1})$, and $V_{2}=(v_{1}^{2}-(s/w)v_{1}+m_{2})$ with $m_{0,1,2}$ and $v_{0,1}$ as in Fig. 2(b), Fig. 2(c),Fig. 2(e), Fig. 2(f), to obtain Fig. 4(a), Fig. 4(b), Fig. 4(c), Fig. 4(d), respectively, and as in Fig. 2(b), Fig. 2(c), Fig. 2(e), Fig. 2(f), to obtain Fig. 4(a), Fig. 4(b),Fig. 4(c), Fig. 4(d), respectively. Analogously to the case of zero modes, the local features of the bounded modes become prominent when the width of the smooth domain wall increases. This is especially visible in Fig. 4(c), where the crossover between oscillations and decaying behavior occurs far away from the center of the domain wall, in correspondence with a point where the quantity $\widetilde{W}(x)$ changes, i.e., where the field $V(x)$ changes its sign.